OpenStudy (anonymous):

find the integral of F.dr ,where F=ie^x +jx^2,and r=i t^2+j t^3, t: [0.2]

OpenStudy (turingtest):

do you still need help with this?

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

then why did you close it? oh well, so \[\vec r=x(t)\hat i+y(t)\hat j\]so now what are x(t) and y(t)?

OpenStudy (turingtest):

\[\int\limits_C\vec F\cdot d\vec r=\int_a^b\vec F(\vec r(t))\cdot\vec r'(t)dt\]

OpenStudy (anonymous):

thanks,,,i got it now

OpenStudy (turingtest):

welcome